Part A Carbonyl fluoride, COF2, is an important intermediate used in the production of fluorine-containing compounds....


Part A
Carbonyl fluoride, COF2, is an important intermediate used in the production of fluorine-containing compounds. For instance, it is used to make the refrigerant carbon tetrafluoride, CF4 via the reaction
2COF2(g)?CO2(g)+CF4(g),    Kc=4.20
If only COF2 is present initially at a concentration of 2.00 M, what concentration of COF2 remains at equilibrium?
Express your answer with the appropriate units.
[COF2] =


Part B
Consider the reaction
CO(g)+NH3(g)?HCONH2(g),    Kc=0.820
If a reaction vessel initially contains only CO and NH3 at concentrations of 1.00 M and 2.00 M, respectively, what will the concentration of HCONH2 be at equilibrium?
Express your answer with the appropriate units.
Answer #1
Concepts and reason

Equilibrium: Equilibrium is a state where the reactant concentration equals the products’ concentration.

or

Equilibrium: Equilibrium is a state where the rate of the forward reaction is equal to the rate of the backward reaction.

Ionic equilibrium: Ionic equilibrium is a state where the equilibrium exists between an ionic compound and its ions.

Fundamentals

Equilibrium state is depicted below.

reactants
equilibrium
products

Equilibrium constant: Equilibrium constant is defined as the ratio of the concentration of products to the concentrations of reactants.

K-[A+][B]
[AB]

(A)

The given equilibrium equation is drawn below.

2COF,(g) 2
CO(g) +CF, (g)

Equilibrium constant expression is given below.

K = [CO][CE]
[COF,]

ICE table for the given equilibrium reaction is drawn below

CO(g)+CF, (g)
2COF,(g)
A
Initial 2.00
Concentration -2x
Equilibrium 2.0-2x

Equilibrium constant value is calculated below.

K =
[x][x]
[2.0-2x]

Given that, Ka = 4.20, the x value is calculated below.

4.20= [x][x]
[2.0-2x]
x=0.80M

concentration at equilibrium is calculated below.

[COF,]=2.00-2x
= 2.00-2 (0.80)
=0.40

(B)

The given equilibrium equation is drawn below.

CO(g)+NH, (g): 2HCONH, (g) [K. =0.820]

Equilibrium constant expression is given below.

K-
[HCONH,]
[CO][NH,]

ICE table for the given equilibrium reaction is drawn below

2HCONH, (g)
Initial
Concentration
Equilibrium
CO(g)+NH,(g)
1.00 2.00
-X
-X
1.0-x 2.00-X
+x

Equilibrium constant value is calculated below.

[1.0-x][2.0-x]

Given that Ka = 0.820, the x value is calculated below.

0.820=
[1.0-x][2.0-x]
x=0.544M

HCONH,
concentration at equilibrium is 0.544M.

Ans: Part A

Concentration of at equilibrium is 0.40M

Part B

HCONH,
concentration at equilibrium is 0.544M